# Boosted Neural Decoders: Achieving Extreme Reliability of LDPC Codes for 6G Networks

### v1

@@ Abstract
Ensuring extremely high reliability is essential for channel coding in 6G networks. The next-generation of ultra-reliable and low-latency communications (xURLLC) scenario within 6G networks requires a frame error rate (FER) below $10^{-9}$. However, low-density parity-check (LDPC) codes, the standard in 5G new radio (NR), encounter a challenge known as the error floor phenomenon, which hinders to achieve such low rates. To tackle this problem, we introduce an innovative solution: boosted neural min-sum (NMS) decoder. This decoder operates identically to conventional NMS decoders, but is trained by novel training methods including: i) boosting learning with uncorrected vectors, ii) block-wise training schedule to address the vanishing gradient issue, iii) dynamic weight sharing to minimize the number of trainable weights, iv) transfer learning to reduce the required sample count, and v) data augmentation to expedite the sampling process. Leveraging these training strategies, the boosted NMS decoder achieves the state-of-the art performance in reducing the error floor as well as superior waterfall performance. Remarkably, we fulfill the 6G xURLLC requirement for 5G LDPC codes without the severe error floor. Additionally, the boosted NMS decoder, once its weights are trained, can perform decoding without additional modules, making it highly practical for immediate application.
@@ I Introduction
Achieving extremely high reliability, characterized by a low frame error rate (FER) of $10^{-9}$, is pivotal for realizing applications in the next generation of ultra-reliable and low-latency communications (xURLLC) within 6G networks [1, 2]. In particular, this level of reliability is crucial for mission-critical applications such as industrial automation, telemedicine, and autonomous driving, serving as key performance indicators (KPIs) [3]. From the perspective of channel coding, it is imperative to develop codes and decoders that achieve the extreme reliability with low signal to noise ratios (SNRs). This high reliability is also required for various systems such as solid-state drive (SSD) storage [4], DNA storage [5], and cryptosystems [6].
However, the low-density parity-check (LDPC) codes employed in the data channel of 5G new radio (NR) suffer from the error floor phenomenon, which hinders the achievement of extremely low FERs at low SNRs [7]. The FER performance curve of LDPC codes typically comprises two distinct regions: the waterfall region, characterized by a sharp decline in FERs, and the error floor region, where the decrease in FERs becomes more gradual. Techniques to improve waterfall performance often inadvertently worsen error floor performance. For example, introducing irregularity into the code structure or employing decoders with weighting (e.g., weighted min-sum: WMS) can improve waterfall performance but degrade error floor performance. Moreover, the implementation of quantization, which is essential for efficient hardware realization, can lead to a severe error floor [8].
@@ I-A Main Contributions
    Boosting learning using uncorrected vectors: We first leverage the boosting learning technique [10] from the machine learning domain. This technique employs a sequential training approach where each classifier concentrates on samples incorrectly classified by its preceding classifiers. By following this approach, we split the whole decoding process into two stages, base and post. The post stage decoder uses specific received vectors that are uncorrected in the base stage as its training samples. If the uncorrected (UC) vectors are accumulated in the error floor region of the base stage, the post stage decoder is specifically trained to correct theses UC vectors causing the error floor.
@@
[…]
@@
    Transfer learning: Boosting learning, unlike conventional methods, requires numerous decoding trials to acquire UC vectors. To mitigate this time-consuming sampling process, transfer learning is employed, where initial weights are set by trained weights on different codes. Starting with these pre-trained weights, optimal weights can be obtained by fine-tuning with few samples. Transfer learning speeds up the sampling process by a factor of $50$.
@@
    Data augmentation: We propose a data augmentation method based on importance sampling to further expedite the sampling process. Similar to the importance sampling technique used for predicting error floor performance [7, 15], we selectively amplify noise at error bit positions of given UC vectors. This method increases the probability of encountering UC vectors. By adopting the transfer learning and data augmentation methods, we are able to accelerate the sampling speed by a factor of $361$.
@@
We refer to the NMS decoder trained through these methods as the boosted NMS decoder. For various practical LDPC codes used in WiMAX [16], WiFi [17], and 5G NR [18], the boosted NMS decoder exhibits notable improvements in error floor performance compared to the WMS decoder [19], threshold attenuated MS (TSMA) decoder [11] and other NMS decoders [9, 20]. The boosted NMS decoder also attains comparable performance to the state-of-the-art post-processing method [12], while requiring only one-third the number of iterations. Importantly, this performance enhancement is achieved solely through modifications to the training process. Consequently, once the training is completed, the boosted NMS decoder can be seamlessly incorporated into the existing hardware architectures designed for the WMS decoder, without requiring additional hardware modifications.
@@
    While our previous work showed the achievement of a FER around $10^{-7}$, this paper advances further by achieving the xURLLC FER target of $10^{-9}$ without severe error floors for 5G LDPC codes.
@@
    Further, this paper provides further analysis on the loss function and the flexibility, and also includes comparisons with recent studies [11, 22, 23], enhancing the understanding of the proposed decoder.
@@ I-B Related Works
Error Floor: The error floor of LDPC codes is known to be caused by problematic small graph structures called trapping sets [7]. Therefore, prior research have focused on developing design algorithms to eliminate trapping sets [24][28] and optimizing decoders to escape trapping sets through perturbations in the decoding process [8, 11][13]. The TAMS decoder [11] applies weighting based on the minimum values for each check node and the post-processing decoder [12] utilizes neural networks to obtain information for managing post-processing; however, they does not employ neural decoders. The study [8] utilized a non-uniform qunatization scheme, which is a complementary approach with ours as we assume the uniform quantization.
@@
Neural Decoders: Neural decoders refer to decoders that perform or aid the decoding task using neural networks. Initially, research focused on replacing decoders with fully connected networks [29][32], and later, model-based neural decoders that use structured neural networks derived from algorithms like belief propagation (BP) and MS were explored [9]. Model-based neural decoders have primarily been applied to high-density codes for alleviating performance degradation caused by structural weaknesses such as cycles [9, 33]. More recently, the application of model-based neural decoders has expanded to include a broader range of code classes such as LDPC [20, 22], generalized LDPC [34], cyclic codes [35], several decoder classes including finite alphabet iterative decoders (FAIDs) [36, 37], and layered decoders [38]. Moreover, various attempts have been made to enhance the performance by leveraging machine learning techniques (e.g., hyper-graph [39], active learning [40], pruning [41], and knowledge distillation [42]).
@@
The comparison with various model-based neural decoders is shown in Table I. A notable point of comparison is the training sample. Most studies have used received vectors simply sampled from the additive white Gaussian noise (AWGN) channel. Recent studies [22, 37] generate noisy codeword vectors with intentionally introduced errors in trapping sets. However, this method requires identifying the trapping sets, which is highly complex and feasible only for short codes. In contrast, the proposed boosting method, which collects training samples through decoding of linear complexity, is applicable to longer codes. Moreover, the distribution of the training samples for the post decoder is identical to that of the received vectors encountered during the actual post decoding process, which makes it a more intuitive and natural method.
@@
In [39], the authors improved waterfall performance by integrating hypernetworks into the standard neural decoder architecture. A transformer architecture was investigated to decode several classes of short codes [23]. They achieved enhanced waterfall performance for short codes at the expense of increased training and decoding costs. Although our proposed training techniques are broadly applicable to these augmented neural decoders, our work primarily focuses on improving error floor performance of long codes with the standard NMS decoder under practical conditions. Instead, we provide a comparative analysis for short codes in the waterfall region, which demonstrates that the proposed methods are also effective for short LDPC codes.
@@ II-B Neural Min-Sum Decoding
Let $m^{(\ell)}_{c\rightarrow v}(m^{(\ell)}_{v\rightarrow c})$ represent the message from CN $c$ to VN $v$ (VN $v$ to CN $c$) for iteration $\ell$. $\mathcal{N}(x)$ denotes the neighbor nodes of the node $x$. The initial messages are set as $m^{(0)}_{v\rightarrow c}=m_{v}^{\rm ch}$, $m^{(0)}_{c\rightarrow v}=0$, where $m_{v}^{\rm ch}$ is the channel log-likelihood ratio (LLR) of VN $v$. For $\ell={1,\ldots,\overline{\ell}}$, the NMS decoding algorithm [9] propagates the following messages
$m^{(\ell)}_{v\rightarrow c}$ $=\overline{w}^{(\ell)}_{v}m_{v}^{\rm ch}+\sum_{c^{\prime}\in\mathcal{N}(v)\setminus c}m^{(\ell-1)}_{c^{\prime}\rightarrow v}$ (1)
$m^{(\ell)}_{c\rightarrow v}$ $={w}^{(\ell)}_{c\rightarrow v}\left(\prod_{v^{\prime}\in\mathcal{N}(c)\setminus v}{\rm sgn}\left(m^{(\ell)}_{v^{\prime}\rightarrow c}\right)\right)\min_{v^{\prime}\in\mathcal{N}(c)\setminus v}|m^{(\ell)}_{v^{\prime}\rightarrow c}|.$ (2)
We call $\overline{w}^{(\ell)}_{v}$ and ${w}^{(\ell)}_{c\rightarrow v}$ the VN weight and the CN weight, respectively. Finally, output LLRs $m_{v}^{\rm o}$ are computed at the last iteration $\overline{\ell}$ by
$m_{v}^{\rm o}=m_{v}^{\rm ch}+\sum_{c\in\mathcal{N}(v)}m^{({\ell})}_{c\rightarrow v}.$
Figure 1: An example of introducing weighting that enables successful decoding.
The quantized NMS decoding algorithm is simply derived by quantizing the values of $m^{(\ell)}_{v\rightarrow c}$, $m^{(\ell)}_{c\rightarrow v}$, and $m_{v}^{\rm ch}$. Quantized decoders are extensively employed in practical applications due to their reduced complexity and hardware-friendly nature. We employ a popular 5-bit uniform quantization setting with a maximum magnitude $7.5$ and a step size $0.5$ to facilitate fair comparison with the existing error floor study [12].
Fig. 1 illustrates an example of MS and NMS decoding, where LLR values $m_{v}^{\rm ch}$, $m^{(\ell)}_{c\rightarrow v}$, and $m_{v}^{\rm o}$ are presented. Initially, two VNs are incorrect, but weighting $m^{(\ell)}_{c\rightarrow v}$ corrects them. Assuming an all-zero codeword, negative values of $m_{v}^{\rm o}$ indicate erroneous VNs. For NMS decoding, $m^{(\ell)}_{c\rightarrow v}$ on black solid edges are weighted by $0.8$, while those on red dotted edges are weighted by $1.3$. The method for assigning weights to $m^{(\ell)}_{c\rightarrow v}$ will be detailed in Subsection III-C. The primary goal of this work is to optimize weights of the NMS decoder to correct error patterns causing the error floor.
@@ II-C Training Weights for the NMS Decoder
Referring to (1) and (2), it is evident that when all weights are set to 1, the NMS decoding reduces to the MS decoding. Alternatively, by assigning $\overline{w}_{v}^{(\ell)}=1,w_{c\rightarrow v}^{(\ell)}=w$ for a single value of $w$, the resulting formula equals to that of the WMS decoder. The diversity of weights in the NMS decoder enhances decoding performance over the WMS and MS decoders. However, assigning distinct weights to every edge, node, and iteration introduces excessive training complexity and memory requirements. To mitigate this, prior researches have implemented weight sharing techniques, assigning the same value to weights having the same attributes. In this study, we primarily employ a protograph weight sharing technique [20], which assigns the identical weight value to the $z$ edges (or nodes) from a proto edge (or node). The weight set is represented by $\{\overline{w}_{v_{p}}^{(\ell)},w_{c_{p}\rightarrow v_{p}}^{(\ell)}\}_{\ell}$ for a proto VN $v_{p}$ and a proto CN ${c_{p}}$, with the total count of $(N+E)\overline{\ell}$. We refer to the weight set as the full weight set. Furthermore, applying temporal weight sharing [9], which unifies weights across all iterations into a single value, results in the weight set $\{\overline{w}_{v_{p}},w_{c_{p}\rightarrow v_{p}}\}$ with a total count of $(N+E)$. On the other hand, utilizing spatial weight sharing [33], where each iteration employs a single VN weight and CN weight, simplify the weight set to a form $\{\overline{w}^{(\ell)},w^{(\ell)}\}_{\ell}$, with the total count adjusting to $2{\overline{\ell}}$.
@@
Fig. 2 depicts a Tanner graph of an LDPC code and the corresponding neural network with ¯$\overline{\ell}=2$. The neural network receives the channel LLR vector $(m_{1}^{\rm ch},\ldots,m_{n}^{\rm ch})$ as input and produces the output LLR vector $(m_{1}^{\rm o},\ldots,m_{n}^{\rm o})$. For each decoding iteration, two hidden layers are allotted, and each layer contains nodes (neurons) whose number is equal to the number of edges in the Tanner graph. The VNs to CNs message in (1) is generated by the odd hidden layers, whereas the even layers pass the CN to VN message in (2). The input layer connected to all odd hidden layers represents the incorporation of the channel LLR in (1). Messages passed from the $2\ell$-th to the $(2\ell+1)$-th hidden layer are weighted by CN weights ${w}^{(\ell)}_{c\rightarrow v}$, and those from the input layer to the $(2\ell+1)$-th hidden layer are weighted by VN weights $\overline{w}^{(\ell)}_{v}$. For the loss function, there have been several loss functions such as the BCE [9], soft bit error rate (BER) [33], and FER loss [37] functions. We implement the FER loss $\frac{1}{2}\left[1-{\rm sgn}\left({\rm min}~{}m_{v}^{\rm o}\right)\right]$ as our target is to minimize the FER in the error floor region.
@@
To train the weights $\overline{w}^{(\ell)}_{v}$ and ${w}^{(\ell)}_{c\rightarrow v}$, we feed training samples into the neural network. Using channel modeling, such as the AWGN channel model, we generate a random received vector of length $n$ along with its corresponding channel LLR vector $(m_{1}^{\rm ch},\ldots,m_{n}^{\rm ch})$ to serve as a training sample. Received vectors are generated assuming the all-zero codeword. An important decision in training is to select the sampling points, or more specifically, the SNR regions for generating received vectors. The SNR range within the waterfall region is typically selected to enhance waterfall performance. Therefore, a straightforward approach for improving error floor performance is to choose the SNR range within the error floor region. However, if training samples are collected from high SNRs targeting the error floor region, the loss function approaches zero, preventing weight updates. Consequently, a more sophisticated technique is required to train the NMS decoder with superior error floor performance.

### v2

@@ Abstract
Ensuring extremely high reliability in channel coding is essential for 6G networks. The next-generation of ultra-reliable and low-latency communications (xURLLC) scenario within 6G networks requires frame error rate (FER) below $10^{-9}$. However, low-density parity-check (LDPC) codes, the standard in 5G new radio (NR), encounter a challenge known as the error floor phenomenon, which hinders to achieve such low rates. To tackle this problem, we introduce an innovative solution: boosted neural min-sum (NMS) decoder. This decoder operates identically to conventional NMS decoders, but is trained by novel training methods including: i) boosting learning with uncorrected vectors, ii) block-wise training schedule to address the vanishing gradient issue, iii) dynamic weight sharing to minimize the number of trainable weights, iv) transfer learning to reduce the required sample count, and v) data augmentation to expedite the sampling process. Leveraging these training strategies, the boosted NMS decoder achieves the state-of-the art performance in reducing the error floor as well as superior waterfall performance. Remarkably, we fulfill the 6G xURLLC requirement for 5G LDPC codes without a severe error floor. Additionally, the boosted NMS decoder, once its weights are trained, can perform decoding without additional modules, making it highly practical for immediate application. The source code is available at https://github.com/ghy1228/LDPC_Error_Floor.
@@ I Introduction
Achieving extremely high reliability, characterized by low frame error rate (FER) of $10^{-9}$, is pivotal for realizing applications in the next generation of ultra-reliable and low-latency communications (xURLLC) within 6G networks [1, 2]. In particular, this level of reliability is crucial for mission-critical applications such as industrial automation, telemedicine, and autonomous driving, serving as key performance indicators (KPIs) [3]. From the perspective of channel coding, it is imperative to develop codes and decoders that achieve extreme reliability with low signal to noise ratios (SNRs). This high reliability is also required for various systems such as solid-state drive (SSD) storage [4], DNA storage [5], and cryptosystems [6].
However, the low-density parity-check (LDPC) codes employed in the data channel of 5G new radio (NR) suffer from the error floor phenomenon, which hinders the achievement of extremely low FER at low SNRs [7]. The FER performance curve of LDPC codes typically comprises two distinct regions: the waterfall region, characterized by a sharp decline in FER, and the error floor region, where the decrease in FER becomes more gradual. Techniques to improve waterfall performance often inadvertently worsen error floor performance. For example, introducing irregularity into the code structure or employing decoders with weighting (e.g., weighted min-sum: WMS) can improve waterfall performance but degrade error floor performance. Moreover, the implementation of quantization, which is essential for efficient hardware realization, can lead to a severe error floor [8].
@@ I-A Main Contributions
    Boosting learning using uncorrected vectors: We first leverage the boosting learning technique [10] from the machine learning domain. This technique involves a sequential training approach for multiple classifiers, where each classifier concentrates on samples incorrectly classified by its preceding classifiers. Following this approach, we split the whole decoding process into two stages: base and post. The post stage decoder is trained on specific received vectors that remain uncorrected in the base stage. If these uncorrected (UC) vectors are collected in the error floor region of the base stage, the post stage decoder is specifically trained to correct these UC vectors causing the error floor.
@@
    Dynamic weight sharing: We introduce a dynamic weight sharing technique that assigns distinct weights to unsatisfied check nodes (UCNs) and satisfied check nodes (SCNs) dynamically during the decoding process, and shares weights within the same iteration. This method significantly reduces the number of trainable weights while preserving performance due to the increased weight diversity from dynamic allocation.
@@
    Transfer learning: Boosting learning, unlike conventional methods, requires numerous decoding trials to acquire UC vectors. To mitigate this time-consuming sampling process, transfer learning is employed, where initial weights are set by trained weights on different codes. Starting with these pre-trained weights, optimal weights can be obtained by fine-tuning with few samples. Transfer learning speeds up the sampling process by tens of times.
@@
    Data augmentation: We propose a data augmentation method based on importance sampling to further expedite the sampling process. Similar to the importance sampling technique used for predicting error floor performance [7, 15], we selectively amplify noise at error bit positions of given UC vectors. This method increases the probability of encountering UC vectors. By adopting the transfer learning and data augmentation methods, we are able to accelerate the sampling speed significantly.
@@
We refer to the NMS decoder trained through these methods as the boosted NMS decoder. For various practical LDPC codes used in WiMAX [16], WiFi [17], and 5G NR [18], the boosted NMS decoder exhibits notable improvements in error floor performance compared to the WMS decoder [19], threshold attenuated MS (TAMS) decoder [11] and other NMS decoders [9, 20]. The boosted NMS decoder also attains comparable performance to the state-of-the-art post-processing method [12], while requiring only one-third the number of iterations. Importantly, this performance enhancement is achieved solely through modifications to the training process. Consequently, once the training is completed, the boosted NMS decoder can be seamlessly incorporated into the existing hardware architectures designed for the WMS decoder, without requiring additional hardware modifications.
@@
    While our previous work showed the achievement of FER around $10^{-7}$, this paper advances further by achieving the xURLLC FER target of $10^{-9}$ without severe error floors for 5G LDPC codes.
@@
    Further, this paper provides an analysis of the loss function and the flexibility, and includes comparisons with recent studies [11, 22, 23], enhancing the understanding of the proposed decoder.
@@ I-B Related Works
Error Floor: Error floors are typically associated with problematic small graph structures called trapping sets [7]. Therefore, prior research has focused on developing design algorithms to eliminate trapping sets [24][28] and optimizing decoders to escape trapping sets through perturbations in the decoding process [8, 11][13]. The TAMS decoder [11] applies weighting based on the minimum message value for each check node and the post-processing decoder [12] utilizes neural networks to obtain information for managing post-processing; however, they do not employ neural decoders. The study [8] utilized a non-uniform quantization scheme, which is a complementary approach with ours as we will assume the uniform quantization.
@@
Neural Decoders: Neural decoders refer to decoders that perform or aid the decoding task using neural networks. Initially, research focused on replacing decoders with fully connected networks [29][32], and subsequently, more advanced architectures, such as recurrent neural network (RNN) [33, 34], autoencoder [35][37], and transformer [23], were explored. Another significant research direction involves model-based neural decoders that employ structured neural networks derived from algorithms like belief propagation (BP) and message passing (MS) [9]. Model-based neural decoders have primarily been applied to high-density codes for alleviating performance degradation caused by structural weaknesses such as cycles [9, 38, 39]. More recently, the application of model-based neural decoders has expanded to include a broader range of code classes such as LDPC [20, 22, 40, 41], generalized LDPC [42], cyclic codes [43], several decoder classes including finite alphabet iterative decoders (FAIDs) [44, 45], layered decoders [46], and soft-output decoders [47]. Moreover, various attempts have been made to enhance the performance by leveraging machine learning techniques (e.g., hyper-graph [48], active learning [49], pruning [50], and knowledge distillation [51]).
@@
The comparison with various model-based neural decoders is shown in Table I. A notable point of comparison is the training sample. Most studies have used received vectors simply sampled from the additive white Gaussian noise (AWGN) channel. Recent studies [22, 45] generate noisy codeword vectors with intentionally introduced errors in trapping sets. However, this method requires identifying the trapping sets, which is highly complex and feasible only for short codes. In contrast, the proposed boosting method, which collects training samples through decoding of linear complexity, is applicable to longer codes. Moreover, the distribution of the training samples for the post decoder is identical to that of the received vectors encountered during the actual post decoding stage, which makes it a more intuitive and natural method.
@@
In [48], the authors improved waterfall performance by integrating hypernetworks into the standard neural decoder architecture. A transformer architecture was investigated to decode several classes of short codes [23]. They achieved enhanced waterfall performance for short codes at the expense of increased training and decoding costs. Although our proposed training techniques are broadly applicable to these augmented neural decoders, our work primarily focuses on improving error floor performance of moderate and long codes with the standard NMS decoder under practical conditions. Instead, we provide a comparative analysis for short codes in the waterfall region, which demonstrates that the proposed methods are also effective for short LDPC codes.
@@ II-B Neural Min-Sum Decoding
Let ${\hat{m}}^{(\ell)}_{c\rightarrow v}~{}(m^{(\ell)}_{v\rightarrow c})$ represent the message from CN $c$ to VN $v$ (VN $v$ to CN $c$) for iteration $\ell$. $\mathcal{N}(x)$ denotes the neighbor nodes of the node $x$. The initial messages are set as ${\hat{m}}^{(0)}_{c\rightarrow v}=0$ and the channel log-likelihood ratio (LLR) of VN $v$ is denoted by $m_{v}^{\rm ch}$. For $\ell={1,\ldots,\overline{\ell}}$, the NMS decoding algorithm [9] propagates the following messages
$m^{(\ell)}_{v\rightarrow c}$ $=\overline{w}^{(\ell)}_{v}m_{v}^{\rm ch}+\sum_{c^{\prime}\in\mathcal{N}(v)\setminus c}{\hat{m}}^{(\ell-1)}_{c^{\prime}\rightarrow v}$ (1)
${\hat{m}}^{(\ell)}_{c\rightarrow v}$ $={w}^{(\ell)}_{c\rightarrow v}\left(\prod_{v^{\prime}\in\mathcal{N}(c)\setminus v}{\rm sgn}\left(m^{(\ell)}_{v^{\prime}\rightarrow c}\right)\right)\min_{v^{\prime}\in\mathcal{N}(c)\setminus v}|m^{(\ell)}_{v^{\prime}\rightarrow c}|.$ (2)
We call $\overline{w}^{(\ell)}_{v}$ and ${w}^{(\ell)}_{c\rightarrow v}$ the channel weight (ChW) and the CN weight (CW), respectively. Finally, output LLRs $m_{v}^{\rm o}$ are computed at the last iteration $\overline{\ell}$ by
$m_{v}^{\rm o}=m_{v}^{\rm ch}+\sum_{c\in\mathcal{N}(v)}{\hat{m}}^{({\ell})}_{c\rightarrow v}.$
Figure 1: An example involves introducing weights that enable successful decoding in a particular case.
The quantized NMS decoding algorithm is simply derived by quantizing the values of $m^{(\ell)}_{v\rightarrow c}$, ${\hat{m}}^{(\ell)}_{c\rightarrow v}$, and $m_{v}^{\rm ch}$. Quantized decoders are extensively employed in practical applications due to their reduced complexity and hardware-friendly nature. We employ a popular 5-bit uniform quantization setting with a maximum magnitude $7.5$ and a step size $\Delta=0.5$ to facilitate fair comparison with the existing error floor study [12].
Fig. 1 illustrates an example of MS and NMS decoding, where LLR values $m_{v}^{\rm ch}$, ${\hat{m}}^{(\ell)}_{c\rightarrow v}$, and $m_{v}^{\rm o}$ are presented. Initially, two VNs are incorrect, but weighting ${\hat{m}}^{(\ell)}_{c\rightarrow v}$ corrects them. Assuming an all-zero codeword, negative values of $m_{v}^{\rm o}$ indicate erroneous VNs. For NMS decoding, ${\hat{m}}^{(\ell)}_{c\rightarrow v}$ on black dotted edges are weighted by $0.8$, while those on red double edges are weighted by $1.3$. This method for assigning weights to ${\hat{m}}^{(\ell)}_{c\rightarrow v}$ will be detailed in Subsection III-C. The primary goal of this work is to optimize the weights of the NMS decoder to correct error patterns causing the error floor.
@@ II-C Training Weights for the NMS Decoder
Referring to (1) and (2), it is evident that when all weights are set to 1, the NMS decoding reduces to the MS decoding. Alternatively, by assigning $\overline{w}_{v}^{(\ell)}=1,w_{c\rightarrow v}^{(\ell)}=w$ for a single value of $w$, the resulting formula equals that of the WMS decoder. The diversity of weights in the NMS decoder enhances decoding performance over the WMS and MS decoders. However, assigning distinct weights to every edge, node, and iteration introduces excessive training complexity and memory requirements. To mitigate this, prior research has implemented weight sharing techniques, assigning the same value to weights having the same attributes. In this study, we primarily employ the protograph weight sharing technique [20], which assigns the identical weight value to the $z$ edges (or nodes) from a proto edge (or node). The weight set is represented by $\{\overline{w}_{v_{p}}^{(\ell)},w_{c_{p}\rightarrow v_{p}}^{(\ell)}\}_{\ell}$ for a proto VN $v_{p}$ and a proto CN ${c_{p}}$, with the total count of $(N+E)\overline{\ell}$. We refer to the weight set as the full weight set. Furthermore, applying temporal weight sharing [9], which unifies weights across all iterations into a single value, results in the weight set $\{\overline{w}_{v_{p}},w_{c_{p}\rightarrow v_{p}}\}$ with a total count of $(N+E)$. On the other hand, utilizing spatial weight sharing [38, 41], where each iteration employs a single ChW and CW, simplifies the weight set to a form $\{\overline{w}^{(\ell)},w^{(\ell)}\}_{\ell}$, with the total count adjusting to $2{\overline{\ell}}$.
@@
Fig. 2 depicts a Tanner graph of an LDPC code and the corresponding neural network with ¯$\overline{\ell}=2$. The neural network receives the channel LLR vector $(m_{1}^{\rm ch},\ldots,m_{n}^{\rm ch})$ as input and produces the output LLR vector $(m_{1}^{\rm o},\ldots,m_{n}^{\rm o})$. For each decoding iteration, two hidden layers are allotted, and each layer contains one node (or neuron) corresponding to each edge in the Tanner graph. The VNs to CNs messages in (1) are generated by the odd hidden layers, whereas the even layers pass the CN to VN messages in (2). The input layer connected to all odd hidden layers represents the incorporation of the channel LLR in (1). Messages passed from the $2\ell$-th to the $(2\ell+1)$-th hidden layer are weighted by CW ${w}^{(\ell)}_{c\rightarrow v}$, and those from the input layer to the $(2\ell-1)$-th hidden layer are weighted by ChW $\overline{w}^{(\ell)}_{v}$. For the loss function, several options exist, such as the BCE [9], soft bit error rate (BER) [38], and FER loss [41, 45] functions. We implement the FER loss $\frac{1}{2}\left[1-{\rm sgn}\left({\rm min}~{}m_{v}^{\rm o}\right)\right]$ as our target is to minimize the FER in the error floor region.
@@
To train the weights $\overline{w}^{(\ell)}_{v}$ and ${w}^{(\ell)}_{c\rightarrow v}$, we feed training samples into the neural network. Using channel modeling, such as the AWGN channel model, we generate a random received vector of length $n$ along with its corresponding channel LLR vector $(m_{1}^{\rm ch},\ldots,m_{n}^{\rm ch})$ to serve as a training sample. Received vectors are generated assuming the all-zero codeword. An important decision in training is to select the sampling points, or more specifically, the SNR region for generating received vectors. The SNR range within the waterfall region is typically selected to enhance waterfall performance. Therefore, a straightforward approach for improving error floor performance is to choose the SNR range within the error floor region. However, if training samples are collected from high SNRs within the error floor region, the loss function approaches zero, preventing weight updates. Consequently, a more sophisticated technique is required to train the NMS decoder with superior error floor performance.